4. Suppose that only Chinese and Japanese firms are producing mobile phones for the Korean market in the short run. The cost function for all Japanese firms is TCJ(q) 100+q2 and for all Chinese firms is TCC(q) = 80 +q2, which firms will produce phones for the Korean market in the long run?
Answer #1

Answer : From the analysis of the average total cost (ATC):

ATC= TC/Q

For Japanese Firm :

TC= 100+q 2

ATC= (100/q )+q

For Chinese firm :

TC =80+q2

ATC= (80/q )+q

Therefore, ATC is minimum in case of Chinese firm .So Chinese are able to produce at lower cost.korean mobile phones are produced by Chinese.